solve equation (x-2)^2+x^2=10
Answer by rapaljer(4671)

(x-2)^2+x^2=10
x^2-4x+4+x^2-10=0
2x^2 -4x-6=0

Divide both sides by 2:
x^2-2x-3=0

Factor the trinomial:
(x-3)(x+1)=0
x=3 or x=-1

Check:
x=3
(3-2)^2+3^2=10
1^2+3^2=10
1+9=10

x=-1
(-1-2)^2+ (-1)^2=10
(-3)^2+(-1)^2=10
9+1=10

It checks!!

For additional help with Solving Quadratic Equations, please seee my own website. Click on my tutor name "rapaljer" anywhere in algebra.com. There you will see the link to my Homepage. Near the top of my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra", and look for "Chapter 2." See "Section 2.09," solving Quadratic Equations. In this chapter, you will find my own non-traditional explanation of Logarithms, complete with examples, exercises, and ALL the answers. Many of these exercises are also solved IN COLOR on my "Math in Living Color" pages that go along with these sections.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus